OpenStudy (anonymous):

Which of the following represents the radical in simplest form? √x^16

OpenStudy (anonymous):

if you'd have \[\LARGE \sqrt{x^2}\] how would it be? :)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

x^16 = ( x^8 )^2 This might help?

OpenStudy (anonymous):

x^8 x^4 x^7√x The question was which of the following, and I forgot to list the following. So here..

OpenStudy (anonymous):

yes but you need to know this formula: \[\Huge \sqrt[a]{x^b}=\left(x\right)^{\frac ba }\] now you have: \[\Huge \sqrt[2]{x^{16}}=x^?\] how would it be?? (just follow the formula)

OpenStudy (anonymous):

@qwertyfreak still confused? :)

OpenStudy (anonymous):

uh yeah..

OpenStudy (anonymous):

where are you confused

OpenStudy (anonymous):

everywhere, I dont know how to do this

OpenStudy (anonymous):

\[\Huge \sqrt[2]{x^{16}}=\left(x\right)^{\frac{16}{2}}=x^? \]